# Classical periodic orbits in extended phase space for spherical harmonic oscillator with spin-orbit coupling

### v1

@@ Abstract
A complete analysis of classical periodic orbits (POs) and their bifurcations was conducted in spherical harmonic oscillator system with spin-orbit coupling. The motion of the spin is explicitly considered using the spin canonical variables derived by semiclassical approximation to the spin coherent state path integral representation. In addition to the diametric and two circular PO families with frozen spin, solutions that bridge two circular POs are found in which orbital motion is coupled to spin precession. In addition, each bridge encounters a secondary bifurcation on the way from one circular PO to the other and generates a new PO, that survives at higher energies while maintaining a constant period. The generic expressions for those POs are obtained explicitly, and all the above peculiar bifurcation scenarios are described fully analytically.
@@ I Introduction
Classical periodic orbits (POs) play central roles in semiclassical quantization for both integrable and nonintegral Hamiltonian systems[1, 2, 3, 4]. Gutzwiller derived a famous trace formula that represents the quantum density of states in terms of the contributions of classical POs[5], based on the semiclassical approximation to the path integral representation using the stationary phase method. The formula works paricularly well for strongly chaotic systems. In another approach, Balian and Bloch derived the semiclassical density of states for infinite well potential systems of arbitrary boundary shapes[6]. It is based on the multiple reflection expansion of the transition amplitude, and can be applied to both integrable and nonintegrable models. For integrable Hamiltonian systems whose energy spectra can be approximately obtained by the Einstein-Brillouin-Keller quantization rule (an extension of the Bohr-Sommerfeld quantization condition to multiply-periodic systems of arbitrary dimensions), Berry and Tabor showed that the oscillating part of the level density is generally represented in the form of the periodic orbit sum[7]. The extension of the Gutzwiller formula to the case of continuous symmetries was conducted by Creagh and Littlejohn[8, 9]. Its extension to bifurcation regions, where the original form of the Gutzwiller formula breaks down, was proposed by Sieber et al. using the uniform approximations[10, 11, 12, 13] and by Magner et al. using the improved stationary-phase method [14, 15]. In general, semiclassical approximation of quantum mechanical level density
$g_{\rm qm}(E)=\sum_{n}\delta(E-E_{n})$
for a conservative Hamiltonian system is given by a trace formula expressed in the form
$g_{\rm cl}(E)=\bar{g}(E)+\sum_{\rm PO}A_{\rm PO}\cos\left(\tfrac{1}{\hbar}S_{\rm PO}-\tfrac{\pi}{2}\mu_{\rm PO}\right).$ (1)
[…]
One of the most significant advantages in the trace formula is that it can explain the origin of the regular oscillations in level distribution, which is generally cannot be understood by purely quantum mechanical concepts. It is especially useful for analyzing gross shell structures in finite quantum systems. It has been applied to nuclear deformation soon after its derivation[16, 17], and then to various quantum systems such as metallic clusters[18, 19, 20] and quantum dots[21]. Periodic-orbit theory (POT) has also been applied to the problems of quantum transport. The oscillation of magnetoresistance in mesocopic conductors, which is related to the electron density of states at the Fermi energy, is successfully explained by the contributing POs to the semiclassical level density[22, 23]. The significance of PO bifurcations as the microscopic origin of the deformed shell structure has been emphasized in [20, 24, 14, 25, 26, 27, 28], and its effect on the magnetoresistance was discussed in [21].
In addition to the orbital degrees of freedom, quantum particles have a spin degree of freedom. The Dirac Hamiltonian for a charged fermion in an electromagnetic field includes the spin-orbit potential, whose effect can be measured as the spin-orbit splitting of the atomic energy levels. In atomic nuclei, strong spin-orbit coupling is inevitable in the mean field of a nucleon to account for the nuclear “magic numbers”, numbers of constituent particles (neutrons or protons) of closed-shell configurations[29, 30, 31]. Thus, the inclusion of spin in semiclassical POT is important for a deeper understanding of nuclear shell structures and related phenomena.
@@
To deal with spin-orbit coupling in semiclassical theory, two different prescriptions have been proposed. The first one is the coupled-channel WKB method[32, 33] in which the multicomponent wave equation is diagonalized using the Wigner-Weyl calculus and is reduced to a set of uncoupled equations corresponding to the individual channels. This method was applied to the problem of the spin-orbit coupling[34] by treating the spin as a channel. This is a kind of adiabatic approximation in which spin is regarded as a slow variable. The semiclassical trace formula is then constructed in terms of the POs in those adiabatic Hamiltonians that correspond to individual spin channels. The weakness of this prescription lies in the occurrence of the so-called mode conversion. Adiabatic approximation cannot be justified in the mode conversion region where two adiabatic Hamiltonians cross with each other and classical dynamics become singular. To deal with this problem, a use of diabatic representation has been proposed, which allows a spin flip at the mode conversion point to remove the singularity[35]. It seems working well in some numerical applications, however, it remains ambiguous and lacks theoretical basis[36].
In another prescription, the spin degree of freedom is explicitly incorporated into the equations of motion (EOM). The classical phase space is extended by introducing a spin canonical variables. The classical dynamics of the spin are derived by a semiclassical approximation of the spin coherent state path integral representation[37]. Theoretically, the semiclassical approximation is justified for large spins, but in practice, it also gives reasonable results for small spins in many cases. Thus, we can expect it to provide a better prescription for a semiclassical treatment of the spin, free from the problems like mode conversions. The compensation is the increased degrees of freedom, which usually complicate classical motions. Pletyukov and Zeitsef derived a trace formula which explicitly consider the spin motion[38]. They also showed that the formula works well for small spins (such as $s=\hbar/2$ for electrons and nucleons) in case Hamiltonians depend linearly on the spin. Thus, the proposed method can be safely applied to the spin-orbit problem in electron and nucleon systems. The application to the 2-dimensional Rashba model to electrons in quantum dots was performed in [36, 39].
In this paper, I investigate the 3-dimensional isotropic harmonic oscillator model with a simple spin-orbit coupling term:
$H=\frac{{\bm{p}}^{2}}{2M}+\frac{M\omega^{2}{\bm{r}}^{2}}{2}-\kappa({\bm{r}}\times{\bm{p}})\cdot{\bm{s}}.$ (2)
This Hamiltonian is physically important as the mean field model for light nuclei, corresponding to the Nilsson model without $l^{2}$ term[40, 31]. Classical dynamics in the same Hamiltonian have been studied numerically in [41]. Quantum energy eigenvalue problem for the Hamiltonian (2) is exactly solved, and the energy eigenvalues for particles with spin $s=\hbar/2$ are expressed as
[…]
where $n,\mathring{l},\mathring{\jmath}$ are radial, spin, orbital angular and total angular momentum quantum numbers, respectively. Each level specified by $\mathring{\jmath}$ has $(2\mathring{\jmath}+1)$-fold degeneracies because the energy eigenvalues are independent on the magnetic quantum number $m_{j}$ ($-\mathring{\jmath}\leq m_{j}\leq\mathring{\jmath}$).
For a system in which the energy eigenvalues are explicitly given in terms of the quantum numbers, Berry-Tabor formula[7] can be applied as the semiclassical density of states, and the following formula has been obtained in Ref. [36] (see also Appendix A)
$g(E)=\bar{g}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (6)
$\begin{split}\delta g_{L}^{(\pm)}(E)&=A_{L}^{(\pm)}\sin\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right)\\ &\quad+B_{L}^{(\pm)}\cos\left(\frac{2\pi LE}{\hbar\omega(1\mp\mathring{\kappa})}-\frac{\pi}{2}L\sigma_{\pm}\right),\end{split}$
$\delta g_{N}^{(0)}(E)=B_{N}^{(0)}\cos\left(\frac{2\pi NE}{\hbar\omega}-\frac{\pi}{2}N\sigma_{0}\right).$
Thus, the terms $\delta g_{L}^{(\pm)}$ and $\delta g_{N}$ can be interpreted as the contributions of POs with the primitive periods
$T_{\pm}=\dfrac{2\pi}{\omega(1\mp\mathring{\kappa})}=\frac{2\pi}{\omega\mp\kappa s}\quad\mbox{and}\quad T_{0}=\dfrac{2\pi}{\omega},$
respectively. In the coupled-channel WKB prescription, two adiabatic Hamiltonians $H_{\pm}=H_{0}\mp\kappa|{\bm{r}}\times{\bm{p}}|s$ are obtained, and circular PO families with periods $T=T_{+}$ and $T_{-}$ are found in the Hamiltonians $H_{+}$ and $H_{-}$, respectively. There is also a diametric PO family with $T=T_{0}$, but it has the orbital angular momentum $\bm{l}={\bm{r}}\times{\bm{p}}=0$ and lies on the mode conversion point $(H_{+}=H_{-})$. Hence, its contribution cannot be evaluated in this prescription. Another problem is the inconsistency in the semiclassical order of the orbit contributions. In general, the amplitude factor of $D$ parameter family is proportional to $\hbar^{-D/2}$. Since both the circular and diametric orbit families have $D=2$ due to the spherical symmetry, the amplitude is expected to be proportional to $\hbar^{-1}$, however, the leading-order term of the amplitudes $A_{L}^{(\pm)}$ in the Berry-Tabor formula (6) is proportional to 2superscriptPlanck-constant-over-$\hbar^{-2}$, which suggests a contribution of 4-parameter family. Then, it is an interesting question if the classical POs in which spin dynamics is explicitly considered in the EOM could help us overcome the above difficulties and obtain a more appropriate quantum-classical correspondence.
@@
This paper is organized as follows. Sec. II recapitulates the basic concepts of the classical counterpart of the spin and its dynamics derived by the spin coherent-state path integral representation. In Sec. III, special types of classical orbits with frozen spin and their properties are discussed. In Sec. IV, generic solutions to the classical EOM with spin precession are derived. Using this solution, periodic orbits are investigated and their bifurcation scenarios with varying energies are clarified. Sec. V is devoted to a summary and perspective.
@@ II Spin canonical variables and the classical equations of motion
The spin is a quantum mechanical quantity analogous to the angular momentum. In quantum mechanics, the spin operator $\hat{s}_{i}$ satisfies the following angular momentum algebra,
@@
where $\varepsilon_{ijk}$ stands for a completely antisymmetric unit tensor (Levi-Civita symbol), and the sum is taken over identical subscripts in each term following the usual Einstein convention. Spin raising and lowering operators $\hat{s}_{\pm}\equiv\hat{s}_{x}\pm i\hat{s}_{y}$ satisfy
@@
where $\mathring{s}$ represents the spin quantum number and $|\mathring{s},m_{s}\rangle$ represents the spin eigenstate satisfying
$\hat{s}^{2}|\mathring{s},m_{s}\rangle=\hbar^{2}\mathring{s}(\mathring{s}+1)|\mathring{s},m_{s}\rangle$
@@
$\zeta$ is an arbitrary complex number, and $|\zeta\rangle$ forms a non-orthogonal, overcomplete set of spin state vectors. The classical counterparts of the spin variables are defined through the expectation values of the spin operators as
@@
$\dot{{\bm{l}}}=\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}}$ (17)
@@
$\dot{{\bm{l}}}+\dot{{\bm{s}}}=0,\quad{\bm{l}}+{\bm{s}}={\bm{j}}=\mbox{const.}$. (18)
Eqs. (16c) and (17) can be rewritten as
$\dot{{\bm{s}}}=-\kappa{\bm{j}}\times{\bm{s}},\quad\dot{{\bm{l}}}=-\kappa{\bm{s}}\times{\bm{l}}=-\kappa{\bm{j}}\times{\bm{l}}$ (19)
@@ III Orbits with frozen spin
$(s_{\perp}=\sqrt{s^{2}-s_{z}^{2}},\quad\omega_{p}=\kappa j)$
@@
$\dot{z}=\frac{p_{z}}{M},\quad\dot{p}_{z}=-M\omega^{2}z$ (21)
@@
$z(t)=z_{0}\sin\omega t,\quad z_{0}=\sqrt{\frac{2E}{M\omega^{2}}}$ (22)
with period $T=2\pi/\omega$. Since the orientation of the spin (the choice of $z$ axis) is arbitrary, the orbit forms a 2-parameter family under given energy $E$.
@@
On the other hand, one has a planar orbit in $xy$ plane by putting $z=0$. To solve the EOM (21) for $x$ and $y$ parts, it is convenient to define the complex variables
@@
$E=\frac{(Mr_{c}\omega)^{2}}{2M}+\frac{M\omega^{2}r_{c}^{2}}{2}-\kappa l_{z}s_{z}=Mr_{c}^{2}\omega(\omega-\kappa s_{z})=l_{z}\omega_{c}.$
@@
The angular frequencies $\omega_{+}$ and $\omega_{-}$ are of the POs whose orbital angular momenta are parallel and anti-parallel to the spin, respectively (see Fig. 2). Each of them forms a 2-parameter family, as well as the diametric family.
@@
In special cases where $\omega_{+}$ and $\omega_{-}$ are commensurable with each other, all planar orbits become periodic. For a spin-orbit parameter satisfying
@@
with integers $n$ and $m$ being relatively prime ($n>|m|\geq 1$), period of the orbit becomes
@@ IV.1 Bifurcations of circular orbits
$E^{\pm}_{k,m}=\frac{\omega_{\pm}(\frac{m}{k}\omega_{\pm}-\kappa s_{z})}{\kappa},\quad s_{z}=\pm s$ (32)
@@ IV.2 General solution to the equations of motion
For a given total angular momentum ${\bm{j}}=(0,0,j)$, EOM for the spin (19) can be solved, and the spin components are explicitly given as functions of time by Eqs. (20). Using this solution, the EOM for the orbital degrees of freedom are expressed as
$\begin{array}[]{rl}\dot{x}&=\dfrac{p_{x}}{M}-\kappa(zs_{y}-ys_{z}),\\[6.0pt] \dot{y}&=\dfrac{p_{y}}{M}+\kappa(zs_{x}-xs_{z}),\\[6.0pt] \dot{z}&=\dfrac{p_{z}}{M}+\kappa(xs_{y}-ys_{x}),\\[6.0pt] \dot{p}_{x}&=-M\omega^{2}x-\kappa(p_{z}s_{y}-p_{y}s_{z})\\ \dot{p}_{y}&=-M\omega^{2}y+\kappa(p_{z}s_{x}-p_{x}s_{z}),\\ \dot{p}_{z}&=-M\omega^{2}z+\kappa(p_{x}s_{y}-p_{y}s_{x})\end{array}$ (39)
@@
EOM (39) are transformed into a compact form as
@@
Because of the relation ${\bm{j}}={\bm{l}}+{\bm{s}}=(0,0,j)$, $x$ and $y$ components of the orbital angular momentum satisfy $l_{x}=-s_{x}$ and $l_{y}=-s_{y}$. By a further transformation $(u,v)=(\xi+i\eta,\xi-i\eta)$ which replaces the trigonometric functions in the coefficients of EOM with exponential functions, one obtains the relations
$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)-\kappa(l_{z}s_{z}-s_{\perp}^{2})$ (47)
$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2})$ (48)
@@
$\begin{array}[]{rl}\dot{u}&=i(\omega-\kappa s_{z})u-\kappa s_{\perp}\zeta e^{-i\omega_{p}t}\\ \dot{v}&=i(\omega+\kappa s_{z})v-\kappa s_{\perp}\zeta e^{i\omega_{p}t}\\ \dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(ue^{i\omega_{p}t}+ve^{-i\omega_{p}t})\end{array}$ (52)
@@
$\begin{array}[]{rl}\dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(U+V)\\ \dot{U}-i\omega_{p}U&=i(\omega-\kappa s_{z})U-\kappa s_{\perp}\zeta\\ \dot{V}+i\omega_{p}V&=i(\omega+\kappa s_{z})V-\kappa s_{\perp}\zeta\end{array}$ (56)
that are combined into an eigenmode equation
@@
where I have used the relation
@@
In the above, $l_{z}\neq 0$ is assumed, which is satisfied by most of the orbits under consideration. (Possibile POs with $l_{z}=0$ are described in Appendix B.) By applying Eqs. (59) and (60) to Eq. (48), one obtains the relation
@@
Nonvanishing $Z_{0}$ leads to $|Z_{+}|=|Z_{-}|$ and thus to improper result $l_{z}=0$. Therefore, $Z_{0}=0$ and one obtains

### v2

@@ Abstract
A complete analysis of classical periodic orbits (POs) and their bifurcations was conducted in spherical harmonic oscillator system with spin-orbit coupling. The motion of the spin is explicitly considered using the spin canonical variables derived by semiclassical approximation to the spin coherent state path-integral representation. In addition to the diametric and two circular PO families with frozen spin, solutions that bridge two circular POs are found in which orbital motion is coupled to spin precession. In addition, each bridge encounters a secondary bifurcation on the way from one circular PO to the other and generates a new PO that survives at higher energies while maintaining a constant period. The generic expressions for those POs are obtained explicitly, and all the above peculiar bifurcation scenarios are described fully analytically.
@@ I Introduction
Harmonic oscillator (HO) system is one of the fundamental systems both in quantum and classical dynamics. Three-dimensional isotropic HO system shows peculiar simplicity due to the dynamical SU(3) symmetry. In quantum mechanics, it has strongly degenerate equidistant energy levels, and in classical mechanics, all the trajectories are periodic. This system is still integrable and exactly solvable after inclusion of spin-orbit coupling given by a simple scalar product of spin and orbital angular momentum vectors. Such Hamiltonian is expressed as
$H=\frac{{\bm{p}}^{2}}{2M}+\frac{M\omega^{2}{\bm{r}}^{2}}{2}-\kappa{\bm{s}}\cdot({\bm{r}}\times{\bm{p}}).$ (1)
Here $M$ is the particle mass, $\omega$ is the oscillator frequency, and $\kappa$ is the spin-orbit coupling strength. This Hamiltonian appears as a simple model for the nuclear mean field[1, 2]. This type of spin-orbit coupling appears in the nonrelativistic approximation to the Dirac equation for spin $1/2$ charged fermions under Coulomb field. In the nuclear mean field, strong spin-orbit coupling is known to exist for both protons and neutrons which comes from a different origin than the case of the Dirac equation. It plays a significant role in explaining the nuclear “magic numbers” in more realistic mean field potentials[3, 4, 2].
In analyzing single-particle shell structures, namely, gross structures of fluctuations in the single-particle level densities, semiclassical periodic-orbit theory (POT) [5, 6, 7, 8, 9, 10] is very useful. It represents the fluctuating part of the level density as the contribution of classical periodic orbits (POs) in the corresponding classical Hamiltonian system. This POT was applied to analyze the gross shell structures in nuclear and cluster systems[11, 12, 13, 14, 15].
@@
If one wants to apply POT to the system with spin degree of freedom, one has to define the classical counterpart of the spin, which is purely quantum mechanical object. To treat spin-orbit coupling in semiclassical POT, the Wentzel-Kramers-Brillouin (WKB) method was first proposed by Littlejohn and Flynn[16] in which the multicomponent wave equation was diagonalized using the Wigner-Weyl calculus and was reduced to a set of uncoupled equations corresponding to the individual spin channels[17, 18]. This is a kind of adiabatic approximation which treats spin as slow variable. However, it suffers a serious problem of mode conversions where adiabatic approximation breaks down and the classical motion becomes singular. A diabatic correction at the mode conversion point has been proposed[19] which seems to be appropriate in some numerical applications, however, it remains ambiguous and lacks theoretical basis[20].
In another prescription, the spin degree of freedom is explicitly incorporated into the equations of motion (EOM). The classical phase space is extended by introducing a spin canonical variables. The classical dynamics of the spin are derived by a semiclassical approximation of the spin coherent state path integral representation[21]. Along with this prescription, the semiclassical trace formula for system with spin degree of freedom was derived[22, 23] in which spin dynamics was explicitly considered in the classical motion. Although the semiclassical approximation was justified only for large spins, the trace formula was found to be applicable to the system of particles with small spins, such as electrons and nucleons, at least for the Hamiltonians which depend linearly on the spin. The Hamiltonian (1) certainly meets this requirement. This finding increases the significance of investigating classical motion, especially the POs, in the extended phase space where spin motion is taken explicitly in the EOM.
Apart from its physical importance as the nuclear mean-field model, the Hamiltonian (1) provides us an interesting and significant example of the integrable spin-orbit-coupled system. The classical motions in the same Hamiltonian were investigated by Rozmej and Arview[24]. The main purpose of their work was to specify the origin of chaotic nature in the nuclear mean-field model. They found that the classical motion in the anisotropic HO system, which is integrable, turns gradually chaotic with increasing spin-orbit coupling strength, and concluded that the spin-orbit coupling can be regarded as one possible source of nuclear chaoticity. In their paper, they also studied the classical motions in the spherical potential, only numerically by the Poincaré plot for two-dimensional orbits in the angular-momentum projected space. The result indicates the integrability of the classical system, but the analytical studies of the orbit have not been attempted.
Classical-quantum correspondence in the Hamiltonian similar to Eq. (1) was considered by Amann et al. in Ref. [20]. They incorporated anisotropy into the HO potential and applied the WKB method. A good quantum-classical correspondence was found in such an anisotropic case. However, in the spherical case, some POs lie in the mode conversion point and the WKB method cannot be applied. They also found several inconsistencies between contribution of classical POs and the exact quantum spectrum. Then they got interest in applying the trace formula in the extended phase space they develped, but actual application has not yet been made. Some details on this will be discussed later in Sec. V.
In this paper, I give the general analytic solution to the EOM for Hamiltonian (1) by explicitly taking account of the spin degree of freedom using the classical spin variables. The EOM are nonlinear in orbital and spin variables, but the spin motion can be exactly solved in the spherical model, and the spin vector is explicitly given by the function of time. Making use of this solution and applying a subtle time-dependent transformation of variables, the system of equations are found to arrive at time-independent homogeneous equations which can be exactly solved. I also construct all the possible POs. With varying energy, those POs exhibit peculiar interesting bifurcation scenarios, and all those bifurcations are described analytically. This kind of model would be not only valuable for the study of nonlinear dynamics in integrable systems but also be suggestive for the study of nonintegrable systems.
@@
The paper is organized as follows. Section II recapitulates the basic concepts of the classical counterpart of the spin and its dynamics derived by the spin coherent-state path-integral representation. In Sec. III, special types of classical orbits with frozen spin and their properties are discussed. In Sec. IV, generic solutions to the classical EOM with spin precession are derived. Using this solution, POs are investigated and their bifurcation scenarios with varying energies are clarified. Section V is devoted to a summary and discussion on the application of semiclassical trace formula.
@@ II Spin canonical variables and the classical equations of motion
Spin is a quantum mechanical quantity analogous to the angular momentum. In quantum mechanics, the spin operator $\hat{s}_{i}$ satisfies the following angular momentum algebra,
@@
where $\varepsilon_{ijk}$ stands for a completely antisymmetric unit tensor (the Levi-Civita symbol), and the sum is taken over identical subscripts in each term following the usual Einstein convention. Spin raising and lowering operators $\hat{s}_{\pm}\equiv\hat{s}_{x}\pm i\hat{s}_{y}$ satisfy
@@
where $\mathring{s}$ represents the spin quantum number [^1] In the following, I will use symbols with circle on top of them as the dimensionless version of the variables without circle. For example, modulus of orbital angular momentum $l=\hbar\mathring{l}$ and spin-orbit parameter $\kappa=\mathring{\kappa}\omega/s$, etc. in Sec. V. and $|\mathring{s},m_{s}\rangle$ represents the spin eigenstate satisfying
$\hat{s}^{2}|\mathring{s},m_{s}\rangle=\hbar^{2}\mathring{s}(\mathring{s}+1)|\mathring{s},m_{s}\rangle,$
@@
Here $\zeta$ is an arbitrary complex number and $|\zeta\rangle$ forms a nonorthogonal, overcomplete set of spin state vectors. The classical counterparts of the spin variables are defined through the expectation values of the spin operators as
@@
$\dot{{\bm{l}}}=\kappa({\bm{r}}\times{\bm{p}})\times{\bm{s}},$ (12)
@@
$\dot{{\bm{l}}}+\dot{{\bm{s}}}=0,\quad{\bm{l}}+{\bm{s}}={\bm{j}}=\mbox{const.}.$. . (13)
Equations (11c) and (12) can be rewritten as
$\dot{{\bm{s}}}=-\kappa{\bm{j}}\times{\bm{s}},\quad\dot{{\bm{l}}}=-\kappa{\bm{s}}\times{\bm{l}}=-\kappa{\bm{j}}\times{\bm{l}},$ (14)
@@ III Orbits with frozen spin
$(s_{\perp}=\sqrt{s^{2}-s_{z}^{2}},\quad\omega_{p}=\kappa j),$
@@
$\dot{z}=\frac{p_{z}}{M},\quad\dot{p}_{z}=-M\omega^{2}z,$ (16)
@@
$z(t)=z_{0}\sin\omega t,\quad z_{0}=\sqrt{\frac{2E}{M\omega^{2}}},$ (17)
with period $T=2\pi/\omega$. Since the orientation of the spin (the choice of $z$ axis) is arbitrary, the orbit forms a two-parameter family under given energy $E$.
@@
Alternatively, putting $z=0$ yields planar orbits in $xy$ plane. To solve the EOM (16) for $x$ and $y$ parts, it is convenient to define the complex variables
@@
$E=\frac{(Mr_{c}\omega)^{2}}{2M}+\frac{M\omega^{2}r_{c}^{2}}{2}-\kappa l_{z}s_{z}=Mr_{c}^{2}\omega(\omega-\kappa s_{z})=l_{z}\omega_{c},$
@@
The angular frequencies $\omega_{+}$ and $\omega_{-}$ are of the POs whose orbital angular momenta are parallel and antiparallel to the spin, respectively (see Fig. 2). Each of them forms a two-parameter family, as well as the diametric family.
@@
In special cases where $\omega_{+}$ and $\omega_{-}$ are commensurable with each other, all planar orbits become perisodic. For a spin-orbit parameter satisfying
@@
with integers $n$ and $m$ being relatively prime ($n>|m|\geq 1$), the period of the orbit becomes
@@ IV.1 Bifurcations of circular orbits
$E^{\pm}_{k,m}=\frac{\omega_{\pm}(\frac{m}{k}\omega_{\pm}-\kappa s_{z})}{\kappa},\quad s_{z}=\pm s.$ (27)
@@ IV.2 General solution to the equations of motion
For a given total angular momentum ${\bm{j}}=(0,0,j)$, the EOM for the spin (14) can be solved, and the spin components are explicitly given as functions of time by Eqs. (15). Using this solution, the EOM for the orbital degrees of freedom are expressed as
$\begin{array}[]{rl}\dot{x}&=\dfrac{p_{x}}{M}-\kappa(zs_{y}-ys_{z}),\\[6.0pt] \dot{y}&=\dfrac{p_{y}}{M}+\kappa(zs_{x}-xs_{z}),\\[6.0pt] \dot{z}&=\dfrac{p_{z}}{M}+\kappa(xs_{y}-ys_{x}),\\[6.0pt] \dot{p}_{x}&=-M\omega^{2}x-\kappa(p_{z}s_{y}-p_{y}s_{z}),\\ \dot{p}_{y}&=-M\omega^{2}y+\kappa(p_{z}s_{x}-p_{x}s_{z}),\\ \dot{p}_{z}&=-M\omega^{2}z+\kappa(p_{x}s_{y}-p_{y}s_{x})\end{array}$ (34)
@@
the EOM (34) are transformed into a compact form as
@@
Because of the relation ${\bm{j}}={\bm{l}}+{\bm{s}}=(0,0,j)$, the $x$ and $y$ components of the orbital angular momentum satisfy $l_{x}=-s_{x}$ and $l_{y}=-s_{y}$. By a further transformation $(u,v)=(\xi+i\eta,\xi-i\eta)$ which replaces the trigonometric functions in the coefficients of EOM with exponential functions, one obtains the relations
$E=\frac{M\omega^{2}}{2}\left(\frac{|u|^{2}+|v|^{2}}{2}+|\zeta|^{2}\right)-\kappa(l_{z}s_{z}-s_{\perp}^{2}),$ (42)
$l_{z}=\frac{M\omega}{4}(|u|^{2}-|v|^{2}),$ (43)
@@
$\begin{array}[]{rl}\dot{u}&=i(\omega-\kappa s_{z})u-\kappa s_{\perp}\zeta e^{-i\omega_{p}t},\\ \dot{v}&=i(\omega+\kappa s_{z})v-\kappa s_{\perp}\zeta e^{i\omega_{p}t},\\ \dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(ue^{i\omega_{p}t}+ve^{-i\omega_{p}t}).\end{array}$ (47)
@@
$\begin{array}[]{rl}\dot{\zeta}&=i\omega\zeta+\tfrac{1}{2}\kappa s_{\perp}(U+V),\\ \dot{U}-i\omega_{p}U&=i(\omega-\kappa s_{z})U-\kappa s_{\perp}\zeta,\\ \dot{V}+i\omega_{p}V&=i(\omega+\kappa s_{z})V-\kappa s_{\perp}\zeta,\end{array}$ (51)
that are combined into a normal mode equation
@@
where I use the relation
@@
In the above, $l_{z}\neq 0$ is assumed, which is satisfied by most of the orbits under consideration. (Possible POs with $l_{z}=0$ are described in the Appendix.) By applying Eqs. (54) and (55) to Eq. (43), one obtains the relation
@@
Nonvanishing $Z_{0}$ leads to $|Z_{+}|=|Z_{-}|$ and thus to the improper result $l_{z}=0$. Therefore, $Z_{0}=0$ and one obtains